OpenStudy (anonymous):

Devon purchased tickets to an air show for 9 adults and 2 children. The total cost was $252. The cost of a childʹs ticket was $6 less than the cost of an adultʹs ticket. Find the price of an adultʹs ticket and a childʹs ticket.

OpenStudy (e.cociuba):

Let C be the cost of a child's ticket and A be the cost of an adult's ticket. 9 adults would spend 9A 2 children would spend 2C If the total was 252, then 9A + 2C = 252 If the cost of a child's ticket is $6 less than the cost of an adult ticket, then C = A - 6 Substituting C = A - 6 into 9A + 2C = 252, we get 9A + 2(A - 6) = 252 9A + 2A - 12 = 252 11A - 12 = 252 11A = 264 A = 24 Substituting A = 22 into C = A - 6, we get C = 24 - 6 = 18 Check 9 adult and 2 child's tickets would cost 9*24 + 2*18 = 216 + 36 = 252 The cost of a child's ticket, 18, is 6 less than the cost of an adult's ticket, 24. So the cost of an adult's ticket is $24 and the cost of a child's ticket is $18.